# The Susceptibility Paradox in Online Social Influence

### v1

@@ Abstract
Understanding susceptibility to online influence is crucial for mitigating the spread of misinformation and protecting vulnerable audiences. This paper investigates susceptibility to influence within social networks, focusing on the differential effects of influence-driven versus spontaneous behaviors on user content adoption. Our analysis reveals that influence-driven adoption exhibits high homophily, indicating that individuals prone to influence often connect with similarly susceptible peers, thereby reinforcing peer influence dynamics. Conversely, spontaneous adoption shows significant but lower homophily. Additionally, we extend the Generalized Friendship Paradox to influence-driven behaviors, demonstrating that users’ friends are generally more susceptible to influence than the users themselves, de facto establishing the notion of Susceptibility Paradox in online social influence. This pattern does not hold for spontaneous behaviors, where friends exhibit fewer spontaneous adoptions. We find that susceptibility to influence can be accurately predicted using friends’ susceptibility alone, while predicting spontaneous adoption requires additional features, such as user metadata. These findings highlight the complex interplay between user engagement and preferences in spontaneous content adoption. Our results provide new insights into social influence mechanisms and offer implications for designing more effective moderation strategies to protect vulnerable audiences.
@@ Introduction
In this paper, we present an empirical framework designed to understand, model, and predict users’ susceptibility to influence on social media platforms. This framework takes into account the networked nature of social influence, with a focus on homophilous influence dynamics, and investigates the Generalized Friendship Paradox, examining how it manifests in influence-driven versus spontaneous behaviors. In detail, we address the following Research Questions (RQs):
@@ 3. RQ3:
By leveraging two large-scale Twitter datasets covering political and public health discussions, our study reveals that susceptibility to influence is homophilous within social networks. We demonstrate that influence-driven behaviors exhibit greater homophily than spontaneous behaviors. Additionally, we found that users’ friends are generally more susceptible to influence and less prone to spontaneous actions than the users themselves, supporting that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous action in friendship networks. These insights inform predictive models, which can predict users’ rate of influence-driven adoptions by leveraging only their friends’ susceptibility to influence. In contrast, spontaneous behaviors are more difficult to predict using solely friends’ influence signals, suggesting a complex interplay of user engagement and preferences in spontaneous actions.
@@
Research also suggests that the GFP could bias local perceptions and individuals’ understanding of societal norms (Alipourfard et al. 2020). Christakis and Fowler (2013) highlight that the spread of behaviors such as smoking, drinking, and exercising is influenced by perceptions of peers’ behaviors. Similarly, susceptibility to adopting information might increase when individuals perceive them as common among their friends, thereby reinforcing the influence of networks on information-sharing behaviors.
@@
Susceptibility to online influence is commonly modeled in probabilistic terms, indicating the likelihood that a user’s behavior and/or beliefs may be influenced by online social interactions (Zhou and Zafarani 2019; Romero, Meeder, and Kleinberg 2011; Goyal, Bonchi, and Lakshmanan 2010). Existing literature consistently highlights two primary components of susceptibility to influence: Exposure to content and the potential subsequent Adoption of that content (Aral and Walker 2012; Hoang and Lim 2016). We adhere to this formulation to construct our Susceptibility Framework, which aims to estimate users’ susceptibility to influence based on the dynamics between exposure and adoption.
@@
On social media, exposure to a piece of content typically refers to a user encountering or viewing the content. Measuring exposure from observational social media data can be challenging, as detailed user behavioral data, such as scrolling and clicking, remain largely inaccessible to academics. Additionally, we don’t have access to the structure of the user follower network and, therefore, whose content each user is exposed to from those whom they follow. The situation is further complicated by algorithmic feed curation, i.e., recommendation algorithms displaying potentially relevant content to a user that does not originate from the users they are following (Guess, Nyhan, and Reifler 2020). Despite these challenges, several studies have approximated exposure by analyzing user interactions (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2022; Sasahara et al. 2021; Ye et al. 2024). The core assumption is that any form of engagement—whether retweeting, quoting, or replying—indicates a user’s potential exposure to another user’s content. These interactions imply that the user has not only viewed but also interacted with the content, thereby increasing the likelihood of subsequent exposures to that user’s content. The rationale is to leverage past interactions as a proxy for predicting future exposure.
@@
In this study, we define adoption as the action of sharing a URL, whether through an original tweet, a retweet, a quoted tweet, or a reply. We focus on URLs because they are commonly embedded in tweets and correspond to specific pieces of information, allowing for straightforward tracking of their diffusion and adoption within a social network. Sharing a URL, whether through original content or interactions, represents a form of information adoption, reflecting a user’s active decision to disseminate specific content. However, it is crucial to note that sharing or adopting a URL does not necessarily imply trust or endorsement of its content. While retweeting is often viewed as an endorsement (Boyd, Golder, and Lotan 2010; Metaxas et al. 2015), other sharing activities can serve to question or critique the content (Hemsley et al. 2018).
Figure 1: An illustration of our susceptibility framework.
@@
Building on our conceptualization of Exposure and Adoption, we develop a framework to estimate user susceptibility to influence. Specifically, by analyzing user activity, we evaluate two crucial dimensions: potential exposure to the content and potential adoption of the content. It is important to acknowledge that our data is ultimately an incomplete rendition of the real world: There may be instances where users have shared or been exposed to content outside of our dataset.
@@
Based on this framework, and in line with findings by Aral and Walker (2012), we categorize adoption into two main types: influence-driven adoption, where the content is shared post-exposure, and spontaneous adoption, where the content is shared without any prior exposure. Recognizing these distinctions is crucial, particularly since adoptions can occur independently of direct influence from others. Motivated by this categorization, we introduce two susceptibility metrics: the Influence-Driven Adoption Rate (IAR) and the Spontaneous Adoption Rate (SAR).
@@
The distributions of the IAR and SAR metrics for both datasets are shown in Figure 6 (see Appendix). The strong negative Spearman correlation between IAR and SAR (Election dataset: $\rho=-0.855$, $\textit{p}<0.001$; Covid dataset: $\rho=-0.857$, $\textit{p}<0.001$) indicates that users who frequently adopt content after being exposed to it (i.e., high IAR) tend to adopt content without prior exposure less frequently (i.e., low SAR), and vice versa. This observation supports the validity of our susceptibility framework and particularly these two metrics, in capturing distinct behaviors on social media—whether users are more influenced by external sources (high IAR) or more prone to independent (i.e., spontaneous) content sharing (high SAR). It is important to note that a high susceptibility level is typically associated with a high IAR and a low SAR. We maintain these two distinct metrics to more accurately reflect the nuances of influence dynamics and user behaviors.
Data Collection
Table 1: Number of users and tweets, along with the number of nodes ($|V|$) and edges ($|E|$) in the user networks $G(V,E)$.
@@
Since our methodology relies on tracking users’ exposure to and adoption of information via URLs embedded in shared tweets, in this study, we only include tweets that contain at least one URL. Detailed statistics on the number of tweets with URLs, the number of engaged users, and the distribution of tweet categories are provided in Table 1.
Identifying “Friends”.
We further refine our dataset to isolate a specific subset of target users. In line with the methodology of Nikolov, Flammini, and Menczer (2021), we establish a minimum inclusion criterion: we study only users who shared at least 10 URLs of any kind during the overall observation period (12 months). This threshold ensures that we concentrate on users with sufficient data to reliably assess their susceptibility. Applying this criterion, our data narrows down to 698,808 users for the Election dataset and 1,047,566 users for the Covid dataset.
We define two users as “friends” if they engage in active reciprocal interactions. This means both users need to actively engage with the other via retweets, quotes, or replies. Although using the followers network could provide a more robust definition of friendship, such data is challenging to obtain using the Twitter API (Martha, Zhao, and Xu 2013). In lieu of the followers data, we use interactions (retweet, quote, and reply) as a proxy for friendship, consistent with previous studies (Jiang, Ren, and Ferrara 2023; Rao, Morstatter, and Lerman 2022). After filtering for users in reciprocal friendships, we have 49,794 and 73,960 users from the Election and Covid datasets, respectively.
Constructing Friendship Networks.
We construct three distinct friendship networks based on the types of reciprocal interactions: Interaction, Retweet, and Mention networks. The Interaction network includes all forms of interactions between users, such as retweets, quotes, and replies. In the Retweet network, connections capture retweets only, whereas the Mention network includes interactions through replies or quotes. We differentiate the Retweet and Mention networks due to their varying social implications and levels of endorsement, a distinction further explored in the “Adoption” section. In Table 1, we present the number of nodes and edges for each network in the two datasets.
@@ Results
Figure 2: The susceptibility metrics (IAR and SAR) exhibit homophily within the social network, as evidenced by the weighted average correlation (Wgt. Avg. Corr.) between users and their friends. All correlation coefficients are significant ($p<0.001$).
Figure 3: Paradox holding probability $p(k,s)$ at varying degrees $k$ using the Interaction network in the Covid dataset.
We first examine whether susceptibility to influence exhibits homophily in a friendship network. To this end, we see if a user’s susceptibility metrics are related to their friends. Following Jiang et al. (2023), we measure homophily by calculating the Pearson correlation between each user’s susceptibility score and the weighted average susceptibility scores of all their friends. This method enables us to consider the frequency of user interactions and accounts for cases where users may have varying levels of influence across different connections. A coefficient of $-1$ suggests that users preferentially connect with others who have opposing susceptibility scores, while a coefficient of $1$ indicates a preference for connections between users with similar susceptibility scores.
The results presented in Figure 2 show that susceptibility is homophilous with respect to both the IAR and SAR scores, as indicated by the statistically significant and positive weighted average correlations. The results are consistent across all types of networks in both datasets. To further confirm the reliability of these findings, we compared our results with those from two null models. The first null model maintains the original network’s degree distribution by randomly selecting and swapping the endpoints of edge pairs. The second null model preserves only the number of edges by randomly reassigning neighbors to nodes at one end of each edge. Both null models result in weighted average correlation measures that are not significant, therefore confirming that our results on the real data cannot be attributed to statistical flukes. Please refer to the Appendix for details.
Our analysis reveals that there is indeed homophily in user susceptibility. Moreover, this homophily, which refers to the tendency for individuals to form connections with others who are similar to them, appears to be more prevalent in influence-driven adoption (IAR) than in spontaneous adoption (SAR). This suggests that within social networks, individuals connected by social ties tend to exhibit similar inclinations toward adopting the behaviors of their friends. This is likely due to the stronger impact of peer influence within close social circles, where interactions are more frequent and intense (De Klepper et al. 2010). As a result, a reinforcing loop emerges: individuals in tight-knit groups are more likely to adopt similar behaviors or attitudes due to the influence of their peers.
@@
Table 2: Empirical results for the Generalized Friendship Paradox in three social networks derived from the Election dataset are presented. For each susceptibility metric $s$ (IAR and SAR), we calculate the Spearman correlation coefficient $\rho_{ks}$ between degree $k$ and $s$ ($p<0.001$ for all $\rho_{ks}$), average paradox holding probability $P$ and the average susceptibility $\langle s\rangle$ of nodes and their friends $\langle s\rangle_{nn}$.
@@
When asking whether the Generalized Friendship Paradox (GFP) holds for susceptibility to influence, we aim to investigate whether individuals’ friends are generally more susceptible than they are. Consistent with Eom and Jo (2014), in this study, we formulate the GFP at both the individual and network levels. At the individual level, a node exhibits GFP if its susceptibility metric is lower than the average of its friends. At the network level, the GFP holds if the average susceptibility metric of all nodes is less than the average susceptibility metric of their respective friends.
Individual-level GFP.
$s_{i}<\frac{\sum_{j\in\mathcal{N}_{i}}s_{j}}{k_{i}},$ (1)
Network-level GFP.
At the network level, the GFP examines the phenomenon across the entire network. It is determined by comparing the average susceptibility metric of every node $\langle s\rangle$ with the average susceptibility metric of all their friends $\langle s\rangle_{nn}$, computed as follows:
$\langle s\rangle_{nn}=\frac{\sum_{i=1}^{N}k_{i}s_{i}}{\sum_{i=1}^{N}k_{i}},$ (2)
where a node $i$ with degree $k_{i}$ is counted as a neighbor $k_{i}$ times. If the average susceptibility metric of a node’s friends exceeds the node’s own average susceptibility metric ($\langle s\rangle<\langle s\rangle_{nn}$), then the GFP holds at the network level.
GFP: Comparing SAR vs. IAR.
Figure 3 depicts the paradox holding probability $p(k,s)$ as a function of degree $k$ and susceptibility metric $s$ using the Interaction network in the Covid dataset. Similar findings from the Election dataset are presented in the Appendix. This three-dimensional representation reveals that, with a constant degree $k$, $p(k,s)$ decreases as $s$ (IAR or SAR) increases, aligning with findings from Eom and Jo (2014). This implies that users with very low susceptibility $s$ tend to connect with those having higher $s$ values, resulting in a $p(k,s)$ close to $1$. Conversely, users with very high susceptibility $s$ values exhibit a $p(k,s)$ approaching 00. Users across this spectrum exhibit a consistent trend, wherein $p(k,s)$ remains close to 1 when $s$ and $k$ are strongly, positively correlated. This effect is particularly evident for IAR, indicating a specific pattern where the paradox holds.
@@
Table 2 presents the empirical results of the individual and network-level GFP across three friendship networks constructed from the Election dataset. We get similar results from the Covid dataset. For each susceptibility metric $s$ (IAR or SAR), we calculate the average paradox holding probability on the individual level $P$, integrating $p(k,s)$ over $k$ and $s$, following Eom and Jo (2014), and the Spearman correlation coefficient between node degree $k$ and susceptibility metric $s$, denoted as $\rho_{ks}$. For the network-level GFP, we compare the node average $\langle s\rangle$ and their friends’ average $\langle s\rangle_{nn}$. We also employ two baseline models for comparison based on the null models described above.
@@
For IAR, the value of $P$ exceeds 0.5 in every friendship network, indicating that the GFP holds at the individual level. Furthermore, the condition $\langle s\rangle<\langle s\rangle_{nn}$ is consistently satisfied across networks, indicating that the GFP also holds at the network level. This implies that, on average, individuals’ friends are more likely to adopt content seen on social media than the individuals themselves. In contrast, for SAR, $P$ consistently falls below 0.5, and $\langle s\rangle$ is always higher than $\langle s\rangle_{nn}$. These observations suggest that the GFP does not hold for spontaneous adoption. In simpler terms, individuals are less inclined to adopt unseen content compared to their friends.
@@
Moreover, it is worth noting that while the GFP is observed for IAR, it is less pronounced compared to the two null models. The discrepancy between individual IAR and that of their friends is smaller in the observed network than in a randomly constructed network. For example, at the individual-level, in the Interaction network of the Election dataset, a random baseline model predicts that in 79.8% of cases, an individual’s friends, on average, would exhibit higher IAR than the individual. However, the actual observed rate in the real-world network is only 54.1%. At the network-level, the differences between $\langle s\rangle$ and $\langle s\rangle_{nn}$ are also smaller in the actual network. This reduction in the discrepancy between a user’s susceptibility and their friends’ average susceptibility is similarly observed for SAR.
@@
These observations corroborate the study by Lerman, Yan, and Wu (2016), which demonstrates that the “majority illusion,” a phenomenon resulting from the friendship paradox, is more intense within networks characterized by heterogeneous degree distributions and disassortative structures. Our observations of real-world social media networks, which exhibit greater homophily compared to random networks, indeed show a reduced paradox holding probability. This implies that the structural attributes of a network can significantly impact the manifestation of the GFP.
@@
Our findings indicate that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous action in friendship networks. This suggests that users’ friends tend to be more susceptible to influence and less spontaneous than the users themselves. Interestingly, the paradox holding probability and the differences in IAR and SAR between users and their friends are less noticeable compared to the random baselines; the less homophilous the network, the larger the discrepancy.
@@
As the GFP provides valuable insights into the relation between an individual’s susceptibility and that of their friends, and both IAR and SAR are homophilous, we investigate the potential to predict a user’s susceptibility based on their friends’ susceptibility levels. Furthermore, we evaluate whether incorporating additional features, such as user and friends’ metadata along with network properties, can enhance prediction accuracy. To accomplish this, we employ both simple linear regression and a random forest model.
@@
Additionally, we also attempt to predict user IAR using friends’ SAR and vice versa; however, these models produce $R^{2}$ values all below 0.05 in both datasets, highlighting that although strongly negative-correlated, IAR and SAR do not adequately predict each other across social networks.
Table 3: Summary of linear regression results for predicting user susceptibility metrics based on friends’ average susceptibility metrics. All regression coefficients reported are significant ($p<0.001$).

### v3

@@ Abstract
Understanding susceptibility to online influence is crucial for mitigating the spread of misinformation and protecting vulnerable audiences. This paper investigates susceptibility to influence within social networks, focusing on the differential effects of influence-driven versus spontaneous behaviors on user content adoption. Our analysis reveals that influence-driven adoption exhibits high homophily, indicating that individuals prone to influence often connect with similarly susceptible peers, thereby reinforcing peer influence dynamics, whereas spontaneous adoption shows significant but lower homophily. Additionally, we extend the Generalized Friendship Paradox to influence-driven behaviors, demonstrating that users’ friends are generally more susceptible to influence than the users themselves, de facto establishing the notion of Susceptibility Paradox in online social influence. This pattern does not hold for spontaneous behaviors, where friends exhibit fewer spontaneous adoptions. We find that susceptibility to influence can be predicted using friends’ susceptibility alone, while predicting spontaneous adoption requires additional features, such as user metadata. These findings highlight the complex interplay between user engagement and characteristics in spontaneous content adoption. Our results provide new insights into social influence mechanisms and offer implications for designing more effective moderation strategies to protect vulnerable audiences.
@@ Introduction
In this paper, we present an empirical framework designed to understand, model, and predict users’ susceptibility to influence on social media platforms. This framework takes into account the networked nature of social influence, with a focus on homophilous influence dynamics, and investigates how the Generalized Friendship Paradox manifests in influence-driven versus spontaneous behaviors, thereby establishing the concept of Susceptibility Paradox. In detail, we address the following Research Questions (RQs):
@@ 3. RQ3:
By leveraging two large-scale Twitter datasets covering political and public health discussions, our study reveals that susceptibility to influence is homophilous within social networks. We demonstrate that influence-driven behaviors exhibit greater homophily than spontaneous behaviors. Additionally, we find that users’ friends are generally more susceptible to influence and less prone to spontaneous actions than the users themselves, supporting that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous action in friendship networks. These insights inform predictive models, which can predict users’ rate of influence-driven adoptions by leveraging only their friends’ susceptibility to influence. In contrast, spontaneous behaviors are more difficult to predict using solely friends’ influence signals, suggesting a complex interplay of user engagement and preferences in spontaneous actions.
@@
Research also suggests that the GFP could bias local perceptions and individuals’ understanding of societal norms (Alipourfard et al. 2020), potentially amplifying or diminishing the spread of information and behaviors. Christakis and Fowler (2013) highlight that the adoption of health-related habits, such as smoking, drinking, and exercising, is influenced by perceptions of peers’ actions. Similarly, an individual’s likelihood of adopting information may be influenced by exposure to friends’ sharing behaviors (such as the frequency and type of content they share), and might increase when content is popular within an ego network. Although this foundational work provides valuable insights into influence dynamics on social media, the relationship between a user’s susceptibility to influence and that of their friends remains underexplored. In this work, our aim is to fill this gap by investigating how network structural properties—specifically homophily and the GFP—shape patterns of susceptibility to influence within social networks.
@@
Susceptibility to online influence is commonly modeled in probabilistic terms, indicating the likelihood that a user’s behavior and/or beliefs may be influenced by online social interactions (Zhou and Zafarani 2019; Romero, Meeder, and Kleinberg 2011; Goyal, Bonchi, and Lakshmanan 2010; Luceri, Braun, and Giordano 2018). The existing literature consistently highlights two primary components of susceptibility to influence: Exposure to content and the potential subsequent Adoption of that content (Aral and Walker 2012; Hoang and Lim 2016). We adhere to this formulation to construct our Susceptibility Framework, which aims to estimate users’ susceptibility to influence based on the dynamics between exposure and adoption.
@@
On social media, exposure to a piece of content typically refers to a user encountering or viewing the content. Measuring exposure from observational social media data can be challenging, as detailed user behavioral data, such as scrolling and clicking, remain largely inaccessible to academics. Additionally, it is increasingly difficult to get access to the structure of the user follower network and, therefore, to whose content each user is exposed to from those whom they follow. The situation is further complicated by algorithmic feed curation, i.e., recommendation algorithms displaying potentially relevant content to a user that does not originate from the users they are following (Guess, Nyhan, and Reifler 2020). Despite these challenges, several studies have approximated exposure by analyzing user interactions (Ferrara and Yang 2015; Rao, Morstatter, and Lerman 2022; Sasahara et al. 2021; Ye et al. 2024). The core assumption is that any form of engagement—whether retweeting, quoting, or replying—indicates a user’s potential exposure to another user’s content. These interactions imply that the user has not only viewed but also interacted with the content, thereby increasing the likelihood of subsequent exposures to that user’s content. The rationale is to leverage past interactions as a proxy to predict future exposure.
@@
In this study, we define adoption as the action of sharing a URL, whether through an original tweet, a retweet, a quoted tweet, or a reply. We focus on URLs because they are commonly embedded in tweets and correspond to specific pieces of information, including multimodal content, such as images and videos, allowing for straightforward tracking of their diffusion and adoption within a social network. Sharing a URL, whether through original content or interactions, represents a form of information adoption, reflecting a user’s active decision to disseminate specific content. However, it is crucial to note that sharing or adopting a URL does not necessarily imply trust or endorsement of its content. Although retweeting is often seen as an endorsement (Boyd, Golder, and Lotan 2010; Metaxas et al. 2015), other sharing activities can serve to question or critique content (Hemsley et al. 2018).
@@
Building on our conceptualization of Exposure and Adoption, we develop a framework to estimate user susceptibility to influence (see Figure 1). Specifically, by analyzing user activity, we evaluated two crucial dimensions: potential exposure to content and potential adoption of content. It is important to acknowledge that our data is ultimately an incomplete rendition of the real world: There may be instances where users have shared or been exposed to content not included in our dataset.
@@
Based on this framework, and in line with the definitions provided by (Aral and Walker 2012), we categorize adoption into two main types: influence-driven adoption, where the content is shared post-exposure, and spontaneous adoption, where the content is shared without any prior exposure. Recognizing these distinctions is crucial, particularly since adoptions can occur independently of direct influence from others. Motivated by this categorization, we introduce two susceptibility metrics: the Influence-Driven Adoption Rate (IAR) and the Spontaneous Adoption Rate (SAR).
@@
The distributions of the IAR and SAR metrics for both datasets are shown in the Appendix (Figure 8). The strong negative Spearman correlation between IAR and SAR (Election dataset: $\rho=-0.855$, $\textit{p}<0.001$; Covid dataset: $\rho=-0.857$, $\textit{p}<0.001$) indicates that users who frequently adopt content after being exposed to it (i.e., high IAR) tend to adopt content without prior exposure less frequently (i.e., low SAR), and vice versa. This observation supports the validity of our susceptibility framework and particularly these two metrics, in capturing distinct behaviors on social media—whether users are more influenced by external sources (high IAR) or more prone to independent (i.e., spontaneous) content sharing (high SAR). It is important to note that a high susceptibility level is typically associated with a high IAR and a low SAR. We maintain these two distinct metrics to more accurately reflect the nuances of influence dynamics and user behaviors.
Data Collection and Curation
@@
Table 1: Number of users and tweets, along with the number of nodes ($|V|$) and edges ($|E|$) in the friendship network $G(V,E)$.
Since our methodology relies on tracking users’ exposed to and adoption of information via URLs embedded in shared tweets, in this study, we only include tweets that contain at least one URL. Detailed statistics on the number of tweets with URLs, the number of engaged users, and the distribution of tweet categories are provided in Table 1.
Selecting “Target Users”.
We further refine our dataset to isolate a specific subset of target users. In line with the methodology of Nikolov, Flammini, and Menczer (2021), we establish a minimum inclusion criterion: we study only users who shared at least 10 URLs of any kind during the overall observation period (12 months). This threshold ensures that we focus on users with sufficient data to reliably assess their susceptibility. Applying this criterion, our data narrows down to 698,808 users for the Election dataset and 1,047,566 users for the Covid dataset.
Figure 2: Differentiation between exposure and friendship. For User X, the unidirectional interaction with User A is modeled as exposure, while the bidirectional interaction with User B is modeled as friendship.
Identifying “Friends” and Constructing Friendship Networks.
Friendship networks are social structures formed by interconnected individuals who are linked through mutual recognition, shared activities, interactions, or sentiments (Feld 1981; Zeggelink, Stokman, and Van De Bunt 1996). We define two target users as “friends” if they engage in active reciprocal interactions. This means both users need to actively engage with the other via retweets, quotes, or replies. Although the followers network could provide a more robust definition of friendship, such data is challenging to obtain using the Twitter API (Martha, Zhao, and Xu 2013). In lieu of the followers data, we use reciprocal interactions (retweet, quote, and reply) as a proxy of friendship, consistent with previous studies (Jiang, Ren, and Ferrara 2023; Rao, Morstatter, and Lerman 2022; Ye et al. 2024). This filtering results in 49,794 and 73,960 users from the Election and Covid datasets, respectively. It is important to note that exposure to a user’s tweet, as defined by unidirectional interactions, does not necessarily indicate friendship with that user. However, the reverse is true: if two users have mutual interactions and are considered friends, they are exposed to each other’s tweets. Figure 2 illustrates the distinction between exposure and friendship.
Figure 3: Homophily of influence‐driven adoption rates (IAR) in the Retweet network of the Election dataset. Nodes are colored by IAR score (blue $=$ low susceptibility, red $=$ high susceptibility), scaled in size by node degree, and connected by edges representing mutual retweet interactions. Panels (a–c) display communities in which connected users share similar IAR scores, illustrating homophily.
We construct three distinct friendship networks based on the types of interactions: Interaction, Retweet, and Mention networks. The Interaction network includes all forms of interaction between users, such as retweets, quotes, and replies. In the Retweet network, connections capture retweets only, whereas the Mention network includes interactions through replies or quotes. We differentiate the Retweet and Mention networks due to their varying levels of social implications and endorsement, a distinction further explored in the “Adoption” section. In Table 1, we present statistical descriptors of each network in the two datasets.
@@ Results
We first examine whether susceptibility to influence exhibits homophily in a friendship network. To this end, we observe whether a user’s susceptibility metrics are related to their friends. Following Jiang et al. (2023), we measure homophily by calculating the Pearson correlation between each user’s susceptibility score and the weighted average susceptibility scores of all their friends. This method enables us to consider the frequency of user interactions and accounts for cases where users may have varying levels of influence across different connections. A coefficient of $-1$ suggests that users preferentially connect with others who have opposing susceptibility scores, while a coefficient of $1$ indicates a preference for connections between users with similar susceptibility scores.
The results presented in Figure 4 show that susceptibility is homophilous with respect to both the IAR and SAR scores, as indicated by the statistically significant and positive weighted average correlations. The results are consistent across all types of networks in both datasets. To further confirm the reliability of these findings, we compared our results with those from two null models. The first null model maintains the original network’s degree distribution by randomly selecting and swapping the endpoints of edge pairs. The second null model preserves only the number of edges by randomly reassigning neighbors to nodes at one end of each edge. Both null models result in weighted average correlation measures that are not significant, therefore confirming that our results on the real data cannot be attributed to statistical flukes. The interested reader can refer to the Appendix for further details.
Figure 4: The susceptibility metrics (IAR and SAR) exhibit homophily within different friendship networks, as evidenced by the weighted average correlation (Wgt. Avg. Corr.) between users and their friends. All correlation coefficients are significant ($p<0.001$).
Figure 5: Paradox holding probability $p(k,s)$ at a varying degree $k$ within the Interaction network of the COVID dataset.
Our analysis reveals that there is indeed homophily in user susceptibility. Moreover, this homophily, which refers to the tendency for individuals to form connections with others who are similar to them, appears to be more prevalent in influence-driven adoption (IAR) than in spontaneous adoption (SAR). Figure 3 visualizes the distribution of IAR scores in the Retweet network of the Election dataset. This suggests that within social networks, individuals connected by social ties tend to exhibit similar inclinations toward adopting the behaviors of their friends. This is likely due to the stronger impact of peer influence within close social circles, where interactions are more frequent and intense (De Klepper et al. 2010). As a result, a reinforcing loop emerges: individuals in tight-knit groups are more likely to adopt similar behaviors or attitudes due to the influence of their peers.
@@
When examining whether the Generalized Friendship Paradox (GFP) applies to susceptibility to influence, we seek to determine whether individuals’ friends are generally more susceptible to influence than they are. Consistent with Eom and Jo (2014), in this study, we formulate the GFP at both the individual and network levels. At the individual-level, a node exhibits GFP if its susceptibility metric is lower than the average of its friends. At the network-level, the GFP holds if the average susceptibility metric of all nodes is less than the average susceptibility metric of their respective friends.
Individual-level GFP.
$s_{i}<\frac{\sum_{j\in\mathcal{N}_{i}}s_{j}}{k_{i}},$ (1)
Network-level GFP.
At the network-level, the GFP examines the phenomenon across the entire network. It is determined by comparing the average susceptibility metric of every node $\langle s\rangle$ with the average susceptibility metric of all their friends $\langle s\rangle_{nn}$, computed as follows:
$\langle s\rangle_{nn}=\frac{\sum_{i=1}^{N}k_{i}s_{i}}{\sum_{i=1}^{N}k_{i}},$ (2)
where a node $i$ with degree $k_{i}$ is counted as a neighbor $k_{i}$ times. If the average susceptibility metric of a node’s friends exceeds the node’s own average susceptibility metric ($\langle s\rangle<\langle s\rangle_{nn}$), then the GFP holds at the network-level.
GFP: Comparing SAR vs. IAR.
Figure 5 depicts the paradox holding probability $p(k,s)$ as a function of degree $k$ and susceptibility metric $s$ using the Interaction network in the Covid dataset. Similar findings from the Election dataset are presented in the Appendix. This three-dimensional representation reveals that, with a constant degree $k$, $p(k,s)$ decreases as $s$ (IAR or SAR) increases, aligning with findings from Eom and Jo (2014). This implies that users with very low susceptibility $s$ tend to connect with those having higher $s$ values, resulting in a $p(k,s)$ close to $1$. Conversely, users with very high susceptibility values exhibit a $p(k,s)$ approaching 00. Users across this spectrum exhibit a consistent trend, wherein $p(k,s)$ remains close to 1 when $s$ and $k$ are strongly, positively correlated. This effect is particularly evident for IAR, indicating a specific pattern where the paradox holds.
@@
Table 2: Empirical results of the Generalized Friendship Paradox (GFP) in the three friendship networks of the Election dataset. For each susceptibility metric $s$ (both IAR and SAR), we calculate the Spearman correlation coefficient $\rho_{ks}$ between degree $k$ and $s$ ($p<0.001$ for all $\rho_{ks}$), average paradox holding probability $P$, and the average susceptibility $\langle s\rangle$ of nodes and their friends $\langle s\rangle_{nn}$.
Table 2 presents the empirical results of the individual- and network-level GFP across three friendship networks constructed from the Election dataset. Similar results are observed in the Covid dataset. For each susceptibility metric $s$ (IAR or SAR), we calculate the average paradox holding probability at the individual-level $P$, integrating $p(k,s)$ over $k$ and $s$, following Eom and Jo (2014), and the Spearman correlation coefficient between node degree $k$ and susceptibility metric $s$, denoted as $\rho_{ks}$. For the network-level GFP, we compare the node average $\langle s\rangle$ and their friends’ average $\langle s\rangle_{nn}$. We also employ two baseline models for comparison based on the null models described above.
@@
For IAR, the value of $P$ exceeds 0.5 in every friendship network, indicating that the GFP holds at the individual-level. Furthermore, the condition $\langle s\rangle<\langle s\rangle_{nn}$ is consistently satisfied across networks, indicating that the GFP also holds at the network-level. This implies that, on average, individuals’ friends are more likely to adopt content they are exposed to than the individuals themselves. In contrast, for SAR, $P$ consistently falls below 0.5, and $\langle s\rangle$ is always higher than $\langle s\rangle_{nn}$. These observations suggest that the GFP does not hold for spontaneous adoption. In simpler terms, individuals are less inclined to adopt unseen content compared to their friends.
@@
Moreover, it is worth noting that while the GFP holds for IAR, it is less pronounced compared to the two null models. The discrepancy between individual IAR and that of their friends is smaller in the observed network than in a randomly constructed network. For example, at the individual-level, in the Interaction network of the Election dataset, a random baseline model predicts that in 79.8% of cases, an individual’s friends, on average, would exhibit higher IAR than the individual. However, the actual observed rate in the real-world network is only 54.1%. At the network-level, the differences between $\langle s\rangle$ and $\langle s\rangle_{nn}$ are also smaller in the actual network. This discrepancy between random baselines and actual networks is similarly observed for SAR.
@@
These observations corroborate the study by Lerman, Yan, and Wu (2016), which demonstrates that the “majority illusion,” a phenomenon resulting from the friendship paradox, is more intense within networks characterized by heterogeneous degree distributions and disassortative structures. Our observations of real-world social media networks, which exhibit greater homophily compared to random baselines, indeed reveal a reduced paradox holding probability. This implies that the structural attributes of a network can significantly impact the manifestation of the GFP.
@@
Our findings indicate that the Generalized Friendship Paradox holds for influence-driven behavior but not for spontaneous actions on social networks. This suggests that users’ friends are generally more susceptible to influence and less prone to spontaneous sharing activities than the users themselves.
@@
As our findings in RQ1 and RQ2 provide valuable insights into the relationship between an individual’s susceptibility and that of their friends, we explore the potential to predict a user’s susceptibility based on the susceptibility levels of their friends. Furthermore, we evaluate whether incorporating additional features, such as user and friend metadata along with network properties, can enhance prediction accuracy. To accomplish this, we employ both simple linear regression and random forest models.
@@
Additionally, we also attempt to predict user IAR using friends’ SAR and vice versa; however, these models produce $R^{2}$ values all below 0.05 in both datasets, highlighting that although strongly negative-correlated, IAR and SAR do not adequately predict each other.
@@
Table 3: Summary of the linear regression results for predicting user susceptibility metrics based on the friends’ average susceptibility metrics. All regression coefficients reported are significant ($p<0.001$).